I tryed working this one it dont seem right
9n+5(n+3)=-(N+13)-2
9N+5N+15+N-11=0
15N+4=0
N=3/4

Answer by rapaljer(4671)

You wrote:
9n+5(n+3)=-(N+13)-2
9N+5N+15+N-11=0

This last step contains an error, and I also see another error in you very last step in your solution. I think in both errors, you tried to do two steps at once. Not usually a good idea, since you don't seem to be handling it without making errors.

I recommend that you remove all parentheses in the first step:
9n + 5n + 15 = -n-13 - 2

The second step is to combine like terms on the left side and also on the right side:
14n + 15 = -n - 15

Next, get all the n terms on one side (the left side) by adding +n to each side:
14n + n + 15 = -n +n - 15
15n + 15 = -15

Next, get all the non-n terms on the other side (right side) by subtracting 15 from each side:
15n + 15 - 15 = -15 -15
15n = -30

Divide both sides by 15:
n= -2

R^2 at SCC